# Harmonic maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$ and the rigidity of the Hopf fibration

### v1

@@ 1. Introduction
$\frac{1}{2}\Delta|B|^{2}=|\nabla B|^{2}\ +\ (\text{terms of order zero in $B$}).$
@@
In our last main result, we consider harmonic maps whose singular values satisfy a suitable identity, without imposing bounds on $B$. In particular, the theorem applies to characterize harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ with constant energy density $|df|^{2}$ or constant 2-dilation $\operatorname{D}_{2}$.
@@
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a non-constant harmonic map whose two largest singular values $\lambda_{1}\geq\lambda_{2}$ satisfy ${W}(\lambda_{1},\lambda_{2})=0,$ for some $C^{1}$-function ${W}:V\subset\mathbb{R}^{2}\to\mathbb{R}$, with $V$ open. If
$y\frac{\partial{W}}{\partial x}+x\frac{\partial{W}}{\partial y}\neq 0,\quad\text{for all}\,\,\,(x,y)\in V\cap W^{-1}(0),$
then $f$ is a Hopf fibration. In particular, if $f$ has constant energy density or constant $2$-dilation, then it must be constant or a Hopf fibration.
@@ 3.2. Harmonic unit vector fields on $\mathbb{S}^{3}$
in the space of unit vector fields on $\mathbb{S}^{3}$, where $d\mu$ denotes the volume element and ${\rm vol}(\mathbb{S}^{3})$ is the volume of $\mathbb{S}^{3}$; see [wiegmink, wood]. Critical points of $E$ for variations by unit vector fields are called harmonic unit vector fields. It is conjectured that the Hopf vector fields are the only harmonic unit vector fields. This conjecture remains open in general, but Fourtzis, Markellos, and Savas-Halilaj [savas2, Theorem B & C] proved it under the additional assumption that $\zeta$ has totally geodesic fibers. To state the result, we first recall the following topological lemma.
@@
Let $\zeta$ be a unit vector field on $\mathbb{S}^{3}$ with compact integral leaves. Then, the quotient space obtained by identifying points on the same fibers of $\zeta$ is a smooth compact surface diffeomorphic to $\mathbb{S}^{2}$, and the quotient map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a smooth fiber bundle (in particular, a submersion).
@@
Let us denote by $\mathbb{M}^{2}$ the space of leaves of the foliation generated by the integral curves of $\zeta$, and denote by $\pi:\mathbb{S}^{3}\to\mathbb{M}^{2}$ the corresponding quotient map, i.e. the map which identifies points of the same integral curve of $\zeta$. Note that $\mathbb{M}^{2}=\pi(\mathbb{S}^{3})$ with the quotient topology is compact and connected. Since $\zeta$ has compact integral fibers, the distribution it generates is regular in the sense of [palais, Chapter 1], thus by [palais, Corollary 5 p.24] the triple
@@
is $\mathbb{S}^{1}$ fiber bundle and $\mathbb{M}^{2}$ is a smooth $2$-dimensional manifold. Moreover, $\mathbb{M}^{2}$ is simply connected, see [hatcher, Theorem 4.41 and Proposition 4.48]. Therefore, $\mathbb{M}^{2}$ is diffeomorphic to the sphere $\mathbb{S}^{2}$. ∎
@@
Let $\zeta\in\mathfrak{X}(\mathbb{S}^{3})$ be a unit vector field whose integral curves are great circles and let $\pi$ be the corresponding quotient map. Then:
@@ 4. Bochner-Weitzenböck formulas
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic map with non-constant singular values $\lambda$ and $\mu$. On the open set where $\lambda$ is different from $\mu$ we have that $\lambda^{2}$ and $\mu^{2}$ are smooth functions, and their Laplacians are given by
@@ 5. Weakly conformal harmonic maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$
The map $g$ is constructed as follows. Given $x\in\mathbb{S}^{2}$, consider the great circle $c_{x}\doteq\pi^{-1}(x).$ Since $f$ and $\pi$ have the same fibers, the circle $c_{x}$ is mapped by $f$ to a point $\tilde{x}=f(c_{x})$. We then define
@@ 6.1. Proof of Theorem A
is not identically zero on $\mathbb{S}^{3}$. Notice that if $v^{2}=\lambda^{2}\mu^{2}=0$ on an open subset of $\mathbb{S}^{3}$, then it is zero everywhere by analyticity and from Lemma 4.3 we deduce that $f$ is constant. Therefore we may assume that the set
@@
is non-empty and open, and the function $w$ is smooth on $U$. Assume that $w$ attains its maximum at a point $p_{0}$. If $p_{0}\in U$, then by Lemma 4.5 we have that
@@ 6.3. Proof of Theorem C
Assume that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a harmonic map whose two largest singular values $(\lambda,\mu)$ satisfy a smooth relation
$W(\lambda,\mu)=0\quad\text{such that}\quad\mu W_{\lambda}+\lambda W_{\mu}\neq 0\quad\text{on $\mathbb{S}^{3}$}.$ (6.7)
Claim 1: Assume that the set
$U=\mathbb{S}^{3}\setminus\{\lambda=\mu\}$
is non-empty. Then $\mu$ is $C^{1}$ on $U$.
Proof of Claim 1. Differentiability may fail on the set
$V=\{\lambda>\mu=0\}\subset\mathbb{S}^{3},$
on which in particular $\lambda$ is smooth; see [nomizu]. Differentiating (6.7) away from $V$ we get that
$W_{\lambda}(\lambda,\mu)\,\nabla\lambda+W_{\mu}(\lambda,\mu)\nabla\mu=0.$
Since $W$ is $C^{1}$ and $\lambda$ and $\mu$ are continuous, the composition $W_{\mu}(\lambda,\mu)$ is continuous and the constraint in (6.7) forces it to be non-zero in an open set $V_{1}$ containing $V$. This implies that, on $V_{1}\setminus V$, we have that
$\nabla\mu=-\frac{W_{\lambda}(\lambda,\mu)}{W_{\mu}(\lambda,\mu)}\,\nabla\lambda.$
@@
Proof of Claim 2. Suppose to the contrary that this is not the case. Then the set $U$ from Claim 1 is non-empty and open, and by Lemma 4.2 the function $\varrho$ is smooth on $U$ and attains its maximum at a point $p_{0}\in U$. Differentiating $\varrho$ and the relation (6.7) at $p_{0}$ we get that
$\begin{cases}W_{\lambda}\,\alpha_{k}(\lambda)+W_{\mu}\,\alpha_{k}(\mu)=0,\\ \lambda\alpha_{k}(\lambda)-\mu\alpha_{k}(\mu)=0,\end{cases}$
@@
It follows that (6.8) must hold everywhere on $\mathbb{S}^{3}$, hence the singular values of $f$ are equal and constant. By [escobales] and (6.7) we conclude that $f$ must be a Hopf fibration.
This immediately implies that, if $\lambda^{2}+\mu^{2}$ or $\lambda\mu$ is equal to a positive constant, then $f$ is a Hopf fibration. If $\lambda^{2}+\mu^{2}$ is identically zero, then $f$ is trivially constant. If $\lambda\mu$ is identically zero, then $\mu$ is identically zero, and it follows from Lemma 4.2 that
$\Delta(\lambda^{2})\geq 4\lambda^{2},$
which forces $\lambda$ to be identically zero as well, therefore $f$ is constant. This concludes the proof.

### v2

@@ 1. Introduction
$\Delta|B|^{2}=2|\nabla B|^{2}\ +\ (\text{terms of order zero in $B$}).$
@@
In our last main result, we consider harmonic maps whose singular values satisfy a suitable identity, without imposing bounds on $B$. In particular, the theorem applies to characterize harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ with constant energy density $|df|^{2}$ or constant 2-dilation $\operatorname{D}_{2}$. To state the result, recall that a function $f:C\subseteq\mathbb{R}^{n}\to\mathbb{R}$ is said to be $C^{1}$ on a subset $C$ if it can be extended to a $C^{1}$-function on an open set containing $C$.
@@
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a non-constant harmonic map whose two largest singular values $\lambda_{1}\geq\lambda_{2}$ satisfy ${W}(\lambda_{1},\lambda_{2})=0,$ for some function $W$ which is continuous on the set
$V=\big\{(x,y)\in\mathbb{R}^{2}\ :\ x\geq y\geq 0\},$
of class $C^{1}$ on $V\cap\{x>y\}\cap W^{-1}(0)$ and there it satisfies
$y\frac{\partial{W}}{\partial x}+x\frac{\partial{W}}{\partial y}\neq 0,\quad\text{for all}\,\,\,(x,y)\in V\cap\{x>y\}\cap W^{-1}(0)$ (1.2)
Then, $f$ is the composition of a Hopf fibration with a holomorphic map $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ (up to conjugation).
Theorem D.
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a non-constant harmonic map with constant energy density or constant $2$-dilation. Then, $f$ is a Hopf fibration.
@@ 3.2. Harmonic unit vector fields on $\mathbb{S}^{3}$
in the space of unit vector fields on $\mathbb{S}^{3}$, where $d\mu$ denotes the volume element and ${\rm vol}(\mathbb{S}^{3})$ is the volume of $\mathbb{S}^{3}$; see [wiegmink, wood]. Critical points of $E$ for variations by unit vector fields are called harmonic unit vector fields. It is conjectured that the Hopf vector fields are the only harmonic unit vector fields. This conjecture remains open in general, but Fourtzis, Markellos, and Savas-Halilaj [savas2, Theorem B & C] proved it under the additional assumption that $\zeta$ has totally geodesic integral leaves. To state the result, we first recall the following topological lemma.
@@
Let $\zeta$ be a unit vector field on $\mathbb{S}^{3}$ with compact integral leaves. Then, the quotient space obtained by identifying points on the same leaves of $\zeta$ is a smooth compact surface diffeomorphic to $\mathbb{S}^{2}$, and the quotient map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a smooth fiber bundle (in particular, a submersion).
@@
Let us denote by $\mathbb{M}^{2}$ the space of leaves of the foliation generated by the integral curves of $\zeta$, and denote by $\pi:\mathbb{S}^{3}\to\mathbb{M}^{2}$ the corresponding quotient map, i.e. the map which identifies points of the same integral curve of $\zeta$. Note that $\mathbb{M}^{2}=\pi(\mathbb{S}^{3})$ with the quotient topology is compact and connected. Since $\zeta$ has compact leaves, the distribution it generates is regular in the sense of [palais, Chapter 1], thus by [palais, Corollary 5 p.24] the triple
@@
is $\mathbb{S}^{1}$ fiber bundle and $\mathbb{M}^{2}$ is a smooth $2$-dimensional manifold. Moreover, by the long exact homotopy sequence of a fiber bundle (see [hatcher, Theorem 4.41 and Proposition 4.48])
$\ldots\to\pi_{1}(\mathbb{S}^{3})\to\pi_{1}(\mathbb{M}^{2})\to\pi_{0}(\mathbb{S}^{1})\to\pi_{0}(\mathbb{S}^{3})\to 0$
we deduce that $\mathbb{M}^{2}$ is simply connected. Therefore, $\mathbb{M}^{2}$ is diffeomorphic to the sphere $\mathbb{S}^{2}$. ∎
@@
Let $\zeta\in\mathfrak{X}(\mathbb{S}^{3})$ be a unit vector field whose integral curves are great circles and let $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ be the corresponding quotient map. Equip the spheres with their standard round metrics. Then:
@@ 4. Bochner-Weitzenböck formulas
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic map. On the open set where $\lambda$ is different from $\mu$, we have that $\lambda^{2}$ and $\mu^{2}$ are smooth functions, and their Laplacians are given by
@@ 5. Weakly conformal harmonic maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$
The map $g$ is constructed as follows. Given $x\in\mathbb{S}^{2}$, consider the great circle $c_{x}\doteq\pi^{-1}(x).$ Since a fiber of $\pi$ is a connected component of a fiber of $f$, the circle $c_{x}$ is mapped by $f$ to a point $\tilde{x}=f(c_{x})$. We then define
@@ 6.1. Proof of Theorem A
is not identically zero on $\mathbb{S}^{3}$. Notice that if $v^{2}=\lambda^{2}\mu^{2}=0$ on an open subset of $\mathbb{S}^{3}$, then it is zero everywhere by analyticity and from Lemma 4.3 we deduce that $f$ is constant. Therefore we may assume that the open set
@@
is non-empty, and observe that $w$ is smooth on $U$. Assume that $w$ attains its maximum at a point $p_{0}$. If $p_{0}\in U$, then by Lemma 4.5 we have that
@@ 6.3. Proof of Theorem C
Denote as usual by $\lambda\geq\mu$ the two largest singular values. Assume that the open set $U=\{\lambda>\mu\}$ is non-empty. By Lemma 4.2, $\lambda^{2}$ and $\mu^{2}$ are smooth on $U$. Moreover, by assumption
$W(\lambda,\mu)=0,\qquad\mu W_{x}(\lambda,\mu)+\lambda W_{y}(\lambda,\mu)\neq 0\quad\text{on $U$}.$ (6.7)
Claim 1: The singular value $\mu$ is $C^{1}$ on $U$.
Proof of Claim 1. Differentiability may fail on the closed analytic subset
$\Sigma=\{\lambda>\mu=0\}\subset U,$
(the zero set of the analytic function $v^{2}$ in (2.4), see Lemma 4.3), on which in particular $\lambda$ is smooth; see [nomizu]. Differentiating (6.7) away from $\Sigma$ we get that
$W_{x}(\lambda,\mu)\,\nabla\lambda+W_{y}(\lambda,\mu)\nabla\mu=0.$
Since $W$ is $C^{1}$ and $\lambda$ and $\mu$ are continuous, the composition $W_{y}(\lambda,\mu)$ is continuous up to $\Sigma$, and the constraint in (6.7) forces it to be non-zero in the entire $U$. This implies the identity
$\nabla\mu=-\frac{W_{x}(\lambda,\mu)}{W_{y}(\lambda,\mu)}\,\nabla\lambda\qquad\text{on }\,U\backslash\Sigma.$
@@
Proof of Claim 2. Suppose to the contrary that this is not the case. Then the open set $U$ from Claim 1 is non-empty, and by Lemma 4.2 the function $\varrho$ is smooth and attains its maximum at a point $p_{0}\in U$. Differentiating $\varrho$ and using the relation (6.7) at $p_{0}$ we get that
$\begin{cases}W_{x}\,\alpha_{k}(\lambda)+W_{y}\,\alpha_{k}(\mu)=0,\\ \lambda\alpha_{k}(\lambda)-\mu\alpha_{k}(\mu)=0,\end{cases}$
@@
By Claim 2, $f$ is a weakly conformal harmonic map. Hence, by Theorem 5.1, the map $f$ is the composition of the Hopf fibration with a conformal map of $\mathbb{S}^{2}$.
6.4. Proof of Theorem D
First assume that the $2$-dilation of $f$ is a constant $c$. Then $f$ satisfies the assumptions in Theorem C with
W(x,y)=xyc.W(x,y)=xy-c.
Hence $f$ is weakly conformal, i.e. $\lambda\equiv\mu$ everywhere on $\mathbb{S}^{3}$. From $W(\lambda,\mu)=0$ we therefore deduce that $\lambda^{2}$ is constant. Escobales’ result [escobales] guarantees that $f$ is a Hopf fibration.
Next assume that the energy density $u$ is a positive constant $c$. From Lemma 4.3 we get
$|B|^{2}=2(v^{2}-u).$ (6.9)
Observe that $\mu$ cannot vanish at any point $p$, otherwise $v$ would vanish at $p$ and thus from (6.9) we get $u=|B|^{2}=0$ at $p$, contradiction. Consequently, $f$ is a submersion. Setting
$W(x,y)=x^{2}+y^{2}-c,$
even though (1.2) is not satisfied on $\{x>y=0\}$, nevertheless
$\mu W_{x}(\lambda,\mu)+\lambda W_{y}(\lambda,\mu)\neq 0\qquad\text{holds everywhere on }\,\mathbb{S}^{3}.$
This suffices to follow the proof of Theorem C and conclude that $f$ is weakly conformal. From $\lambda\equiv\mu$ and $u=\lambda^{2}+\mu^{2}$ constant we deduce that $\lambda$ and $\mu$ are constant. The conclusion follows by Escobales’ result.